# Junhao Chen* 1122    Xiang Li* 33    Xiaojun Ye 1144    Chao Li 55 Zhaoxin Fan 66

### v1

@@ Abstract
In this paper, we pursue a novel 3D AIGC setting: generating 3D content from IDEAs. The definition of an IDEA is the composition of multimodal inputs including text, image, and 3D models. To our knowledge, this challenging and appealing 3D AIGC setting has not been studied before. We propose the novel framework called Idea-2-3D to achieve this goal, which consists of three agents based upon large multimodel models (LMMs) and several existing algorithmic tools for them to invoke. Specifically, these three LMM-based agents are prompted to do the jobs of prompt generation, model selection and feedback reflection. They work in a cycle that involves both mutual collaboration and criticism. Note that this cycle is done in a fully automatic manner, without any human intervention. The framework then outputs a text prompt to generate 3D models that well align with input IDEAs. We show impressive 3D AIGC results that are beyond any previous methods can achieve. For quantitative comparisons, we construct caption-based baselines using a whole bunch of state-of-the-art 3D AIGC models and demonstrate Idea-2-3D out-performs significantly. In 94.2% of cases, Idea-2-3D meets users’ requirements, marking a degree of match between IDEA and 3D models that is 2.3 times higher than baselines. Moreover, in 93.5% of the cases, users agreed that Idea-2-3D was better than baselines. Codes, data and models will made publicly available.
Keywords:
LMM agents 3D AIGC Automated 3D Design
11footnotetext: * Indicates Equal Contribution. Indicates Corresponding Author.
Figure 1: The Idea-2-3D framework synergizes the capabilities of the Large Multimodal Model (LMM), Text-to-Image (T-2-I), and Image-to-3D (I-2-3D) models to transform complex multimodal input (one sentence and two images in this case) IDEAs into tangible 3D models. This process begins with the user articulating high-level 3D design requirements (IDEA). Following this, the LMM generates textual prompts (Prompt Generation) that are then converted into 3D models. These models are evaluated through a Multiview Image Generation and Evaluation process, leading to the Selection of an Optimal 3D Model. Subsequently, the T-2-I prompt is refined (Feedback Generation) using insights from the GPT-4V. Additionally, an integrated memory module (see Sec. 3.7), while not depicted here, meticulously records each iteration, facilitating a multimodal, iterative self-refinement cycle within the framework. Note that this procedure is fully automatic without any human intervention.
@@ 1 Introduction
Recently the success of 2D AIGC foundataion models [40] has been translated to the 3D domain [20, 22, 33]. However, state-of-the-art models take RGB images or text prompts as inputs, which still fails to match the (wildest) creativity of humanity. Arguably, the nature of creativity is connecting (seemingly unrelated) dots that share intrinsic harmony. So we propose a novel 3D AIGC setting in which all prior arts fail: generating textured 3D models from IDEAs. The formal definition of an IDEA is an interleaved sequence of multi-modal inputs, covering modalities like text, images and 3D models. We show a typical IDEA in Fig. 1, which contains one sentence and two images. This kind of IDEAs come into our minds now and then in the daily life: Someone takes a doughnut into the room and looks at her pet rabbit, imaging what it may look like if the rabbit is eating the doughnut using front paws. This is the moment that creativity happens but as far as we know, no existing 3D AIGC foundation models can take this kind of IDEAs as input. More different types of IDEAs are shown in Fig.2.
@@
Existing methods in text-based 3D model generation [50, 33, 17, 49, 25, 48, 20, 34, 22], known as Text-to-3D (T-2-3D), have made progress in certain aspects such as fidelity, but they still face substantial challenges, particularly when dealing with complex and abstract interleaved multimodal inputs (IDEAs). A viable solution that adapts existing T-2-3D methods for IDEA inputs is to turn IDEAs into natural language descriptions. However, this solution can be time-consuming and demands a specific level of expertise for human users.
@@
Our proposal is to use a multi-agent collaboration framework. LLM (Large Language Model) agent systems [54, 38, 46, 39] have already demonstrated remarkable effectiveness in solving complex natural language processing tasks, suggesting their potential application in the T-2-3D domain. There are already some recent successful methods that leverage LLM agents for computer vision applications, like [10, 45, 51]. They exploit the generic methodology of prompting LLM agents to write codes and invoke existing computer vision functions. We inherit this methodology but exploit LMMs (Large Multimodal Models) as agents because the visual inputs are critical in understanding IDEAs.
@@
However, designing a LMM agent system to generate 3D models from IDEAs is not straightforward and presents its own set of challenges, especially effective integration and understanding of multimodal inputs. To tackle these challenges, we propose Idea-2-3D, a framework that employs three different agents based upon the powerful LMM GPT-4V(ision) [26], for iterative self-improvement in automated 3D design and generation. Specifically, as shown in Fig. 1, Idea-2-3D consists of three LMM agents (green boxes indexed by 1,5,6) acting in the roles of prompt generation, model selection and feedback reflection.
@@
As inspired by Idea2img [55], Idea-2-3D combines the capabilities of LMM agents based upon GPT-4V and other multimodal algorithmic modules (purple boxes indexed by 2,3 and yellow box indexed by 4 in Fig. 1) to generate textual prompts from interleaved user inputs (IDEAs), which are then converted into 3D models. This process involves iterative refinement, utilizing a memory module to record each iteration and support continuous improvement. As shown in Fig. 2, Idea-2-3D can generate high-quality 3D models that well align input IDEAs in a fully automatic manner while caption-based baselines constructed from prior T-2-3D models can hardly generate meaningful results.
Our qualitative comparisons and quantitative experiments demonstrate the effectiveness of Idea-2-3D, especially in handling complex and challenging IDEA inputs, where users perceive that 94.2% of the cases are satisfying IDEA requirements. This is also more than twice as good as existing methods (including closed-source commercial methods) at satisfying IDEA requirements.
To summarize, we make the following contributions:
@@
  • Addressing the novel task of fully automated 3D AIGC by transforming high-level, abstract user IDEAs into concrete 3D models, for the first time.
@@
  • Going further than existing LLM agent systems, we demonstrating the effectiveness of agents based upon LMMs represented by GPT-4V in improving, evaluating, and validating multimodal content for 3D model generation.
  • Introducing Idea-2-3D, which leverages T-2-3D model tools to handle interleaved multimodal inputs (IDEAs) and generate high-quality 3D models.
  • Presenting an evaluation set of challenging multimodal IDEAs and confirming Idea-2-3D’s effectiveness through a comprehensive user preference study.
Figure 2: Overview of 3D models generated from various types of multimodal IDEA inputs supported by Idea-2-3D. The light red box on the left is the user input IDEA containing text, images and 3D models. In the center are the baseline results generated directly from the same T-2-I model with caption-based T-2-I prompt (see  LABEL:Compared_Model_Variants). The model on the right is the result generated by iteratively self-refining the T-2-I prompts with Idea-2-3D. Comparison with more existing methods is shown in  LABEL:fig:comp.
@@ 2 Related Works
Self-refining Agents. Our research builds on the self-refinement capability of large language models (LLMs). Recent studies show that LLMs are effective at self-refinement, when exploring unknown environments and tasks [54, 24, 32, 43]. For example, projects such as Self-refine [24] and Reflexion [43] utilize LLMs to iteratively critique their outputs and use feedback to improve prediction.
@@
However, these approaches mainly excel in tasks dealing with natural language descriptions [44]. In contrast, our Idea-2-3D project employs an iterative self-refinement system based on LMM in a multimodal environment, especially for interleaved inputs of text, images, and 3D models (IDEAs), in a different way other than the traditional approach focusing solely on natural language inputs.
Large Multimodal Model. Building on the Large Language Model (LLM), the development of the Large Multimodal Model (LMM) marks an important evolution from unimodal to multimodal processing capabilities. Initial LLMs, such as the GPT family, focused on the generation of textual data, demonstrating superior capabilities in understanding and creating natural language [3, 36].
@@
As the field evolves, the processing capabilities of the models expanded from pure text to include multimodal data including images, audio, and video [37, 35]. For example, CLIP [35] demonstrates cross-modal understanding between text and image, and LLaVA [19] demonstrates multimodal understanding and generation of mixed text, image, and video inputs. Further, projects such as Uni-3D [57] and SDFusion [5] extend this concept to the 3D design domain, enabling good 3D model understanding, generation and reconstruction. LMM has also enabled open-set scene understanding in various settings [47, 15, 16, 21, 11].
Despite the progress made in the field of 3D understanding and generation, current LMMs can still only handle input content at the text and image level, and struggle to handle multimodal high-level inputs containing text, images, and 3D models (IDEAs in our case). In contrast, the Idea-2-3D project takes a much larger step forward in multimodal 3D model generation. Our system is capable of processing not only single-modal inputs, but also composite multimodal inputs containing text, images, and 3D models at the same time.
@@
Extensions of T-2-3D Models. There is already a large literature extending seminal T-2-3D models, including variants enabling T-2-3D models to better follow user prompts [1, 4, 6] , refine keywords in T-2-3D prompts for better visual quality [8], support for additional image inputs for image processing [2, 12], support for additional 3D model inputs for 3D model processing [5], 3D style migration [31, 23, 41], visual concept customization [52, 13], and more. Going even further, Idea-2-3D offers users a more natural way to design and create the 3D content they want. Similar to Visual ChatGPT [53] , which extends ChatGPT’s [30] ability to understand and generate 2D images, Idea-2-3D is designed to provide a more unified and broadly applicable framework for automated 3D model design and generation. Idea-2-3D extends the multimodal input and 3D model understanding and generation capabilities of the LMM [26] using the T-2-I model [42] and I-2-3D model [37], as well as the multimodal input capabilities of the T-2-3D model [50, 33, 17, 49, 25, 48, 34, 20, 22] and other standard off-the-shelf algorithmic modules like multiview rendering.
Figure 3: Overview of the framework of Idea-2-3D, which employs LMM agents to unleash the T-2-3D model’s potential through iterative self-refinement to provide better T-2-3D prompts for the input user IDEA. Green rounded rectangles indicate steps completed by GPT-4V [26] (a.k.a. LMM agents). Purple rounded rectangles indicate T-2-3D modules, including T-2-I models and I-2-3D models. The yellow rounded rectangle indicates the off-the-shelf 3D model multi-view generation algorithm. The blue color indicates the memory module, which saves all the feedback from previous rounds, the best 3D model, and the best text prompt. Note that this cycle is fully automatically executed by LMM agents, without any human intervention.
3 Idea-2-3D Framework
@@
The Idea-2-3D framework represents a novel approach to generate detailed 3D models from high-level, abstract multimodal inputs (IDEAs), shown in Fig. 3. It integrates three LMM-based agents and several off-the-shelf tools for agents to invoke. Specifically, three agents are responsible for prompt generation (green box indexed by 1 in Fig. 3), model selection (green box indexed by 5 in Fig. 3) and feedback generation (green box indexed by 6 in Fig. 3). Two foundation models for T-2-I (purple box indexed by 2 in Fig. 3) and I-2-3D (purple box indexed by 3 in Fig. 3) are exploited together to turn natural language prompts into textured 3D models. As shown in Table. 1, various foundation model variants are evaluated for comprehensiveness. A unique memory module enhances the system, retaining insights from previous iterations to optimize future outputs.
@@ 3.2 Initial Prompt Generation
Here, $p_{\rm gen}$ represents a prompt facilitating the generation of T-2-3D prompts [^2] Note here $p_{\rm gen}$ and P are prompts for LMMs and T-2-3D models, respectively.. This enables Idea-2-3D to extract and interpret not only direct descriptions (e.g. $X^{\prime}$ with only $T$ is also acceptable) but also high-level concepts and intermixed modalities within the IDEA, such as images and 3D models.
@@
Idea-2-3D transforms the set of N text prompts $\{P_{0},P_{1},\ldots,P_{N-1}\}$ into an equivalent set of N 3D models $\{D_{0},D_{1},\ldots,D_{N-1}\}$ utilizing T-2-3D models. As outlined in Sec. 3.2, T-2-3D encompasses a two-step conversion process: initial Text-to-Image (T-2-I) generation followed by Image-to-3D (I-2-3D) generation. To improve 3D model creation quality, we employ a background removal module [7] on T-2-I outputs before I-2-3D processing. The T-2-3D function is :

### v2

# Idea23D: Collaborative LMM Agents Enable 3D Model Generation from Interleaved Multimodal Inputs

@@ Abstract
With the success of 2D diffusion models, 2D AIGC content has already transformed our lives. Recently, this success has been extended to 3D AIGC, with state-of-the-art methods generating textured 3D models from single images or text. However, we argue that current 3D AIGC methods still don’t fully unleash human creativity. We often imagine 3D content made from multimodal inputs, such as what it would look like if my pet bunny were eating a doughnut on the table. In this paper, we explore a novel 3D AIGC approach: generating 3D content from IDEAs. An IDEA is a multimodal input composed of text, image, and 3D models. To our knowledge, this challenging and exciting 3D AIGC setting has not been studied before. We propose the new framework Idea23D, which combines three agents based on large multimodal models (LMMs) and existing algorithmic tools. These three LMM-based agents are tasked with prompt generation, model selection, and feedback reflection. They collaborate and critique each other in a fully automated loop, without human intervention. The framework then generates a text prompt to create 3D models that align closely with the input IDEAs. We demonstrate impressive 3D AIGC results that surpass previous methods. To comprehensively assess the 3D AIGC capabilities of Idea23D, we introduce the Eval3DAIGC-198 dataset, containing 198 multimodal inputs for 3D generation tasks. This dataset evaluates the alignment between generated 3D content and input IDEAs. Our user study and quantitative results show that Idea23D significantly improves the success rate and accuracy of 3D generation, with excellent compatibility across various LMM, Text-to-Image, and Image-to-3D models. Code and dataset are available at https://idea23d.github.io/.
Idea23D: Collaborative LMM Agents Enable 3D Model Generation from Interleaved Multimodal Inputs
Junhao Chen1,2*, Xiang Li3*, Xiaojun Ye4, Chao Li5, Zhaoxin Fan6, Hao Zhao1 1Institute for AI Industry Research (AIR), Tsinghua University, 2Tsinghua Shenzhen International Graduate School, Tsinghua University, 3School of Software and Microelectronics, Peking University, 4College of Computer Science, Zhejiang University, 5College of Computer Science and Technology, Harbin Engineering University, 6Beijing Advanced Innovation Center for Future Blockchain and Privacy Computing, School of Artificial Intelligence, Beihang University
footnotetext: * Indicates Equal Contribution. Indicates Corresponding Author, email to zhaohao@air.tsinghua.edu.cn and zhaoxinf@buaa.edu.cn
{strip}[Uncaptioned image]
Figure 1: The Idea23D framework synergizes the capabilities of the Large Multimodal Model (LMM), Text-to-Image (T-2-I), and Image-to-3D (I-2-3D) models to transform complex multimodal input IDEAs into tangible 3D models. This process begins with the user articulating high-level 3D design requirements (IDEA). Following this, the LMM generates textual prompts (Prompt Generation) that are then converted into 3D models. These models are evaluated through a Multiview Image Generation and Evaluation process, leading to the Selection of an Optimal 3D Model. Subsequently, the T-2-I prompt is refined (Feedback Generation) using insights from the LMM. Additionally, an integrated memory module (see Sec. 3.7), meticulously records each iteration, facilitating a multimodal, iterative self-refinement cycle within the framework. Note that this procedure is fully automatic without any human intervention.
@@ 1 Introduction
Recently the success of 2D AIGC foundation models Rombach et al. (2022); Podell et al. (2023); Zhang et al. (2023c); Shi et al. (2023); Liu et al. (2023f); OpenAI (2023b, c); Yang et al. (2023b); Chen et al. (2023b); Labs (2024) has been translated to the 3D domain Liu et al. (2023d); Long et al. (2023); Poole et al. (2022); Lin et al. (2023); Voleti et al. (2024); Xu et al. (2024b); Tochilkin et al. (2024); Yang et al. (2024). However, state-of-the-art models take RGB images or text prompts as inputs, which still fails to match the (wildest) creativity of humanity. Arguably, the nature of creativity is connecting (seemingly unrelated) dots that share intrinsic harmony. So we propose a novel 3D AIGC setting in which all prior arts fail: generating textured 3D models from IDEAs. The formal definition of an IDEA is an interleaved sequence of multi-modal inputs, covering modalities like text, images and 3D models. We show some typical IDEAs in Fig. 1, which contain a text prompt and images or a 3D model. We use rendered images to represent the 3D model in Fig. 1. This kind of IDEAs come into our minds now and then in the daily life: Someone takes a doughnut into the room and looks at her pet rabbit, imaging what it may look like if the rabbit is eating the doughnut using front paws. This is the moment that creativity happens but as far as we know, no existing 3D AIGC foundation models can take this kind of IDEAs as input. More different types of IDEAs are shown in Fig. 3 and Fig. 8.
@@
Existing methods in text-based 3D model generation Wang et al. (2024b); Poole et al. (2022); Lin et al. (2023); Liu et al. (2023d); Qian et al. (2023); Yang et al. (2024), known as Text-to-3D (T-2-3D), have made progress in certain aspects such as fidelity, but they still face substantial challenges, particularly when dealing with complex and abstract interleaved multimodal inputs (IDEAs). A potential solution for adapting existing T-2-3D methods to handle IDEA inputs is converting the images and 3D models in the IDEAs into natural language descriptions. However, this approach is time-consuming and requires a certain level of expertise from the user.
@@
Our proposal is to use a multi-agent collaboration framework. LLM (Large Language Model) agent systems Xi et al. (2023); Reworkd (2023); Team (2023); Richards (2023); Gong et al. (2023); Liu et al. (2023h); Gu et al. (2023b); Liu et al. (2023a); Chen et al. (2023c); Gou et al. (2023) have already demonstrated remarkable effectiveness in solving complex natural language processing tasks, suggesting their potential application in the T-2-3D domain. There are already some recent successful methods that leverage LLM agents for computer vision applications Gupta and Kembhavi (2023); Surís et al. (2023); Wei et al. (2024); Yang et al. (2023a); Huang et al. (2023); Yang et al. (2023b); Ye et al. (2024). They exploit the generic methodology of prompting LLM agents to write codes and invoke existing computer vision functions. We inherit this methodology but exploit LMMs (Large Multimodal Models) as agents because the visual inputs are critical in understanding IDEAs.
@@
However, designing a LMM agent system to generate 3D models from IDEAs is not straightforward and presents its own set of challenges, especially effective integration and understanding of multimodal inputs. To tackle these challenges, we propose Idea23D, a framework that employs three different agents based upon the powerful LMM, for iterative self-improvement in automated 3D design and generation. Specifically, as shown in Fig. 1, Idea23D consists of three LMM agents (green boxes indexed by 1,5,6) acting in the roles of prompt generation, model selection and feedback reflection.
@@
Idea23D combines the capabilities of LMM agents and other multimodal algorithmic modules (purple boxes indexed by 2,3 and yellow box indexed by 4 in Fig. 1) to generate textual prompts from interleaved user inputs (IDEAs), which are then converted into 3D models. This process involves iterative refinement, utilizing a memory module to record each iteration and support continuous improvement. As shown in Fig. 3 and Fig. 8, Idea23D can generate high-quality 3D models that well align input IDEAs in a fully automatic manner while caption-based baselines constructed from prior T-2-3D models can hardly generate meaningful results.
Our qualitative comparisons and quantitative experiments demonstrate the effectiveness of Idea23D, especially in handling complex and challenging IDEA inputs. The contributions of this paper are as follows.
  1. (1)
@@
    Idea23D is the first work to achieve the transformation of high-level, abstract user IDEAs (multimodal interleaved inputs) into concrete 3D models, realizing a fully automated 3D AIGC task.
  2. (2)
@@
    Surpassing the capabilities of existing LLM agent systems in 3D AIGC, Idea23D demonstrates the effectiveness of LMM-based agents in improving, evaluating, and validating multimodal content for 3D model generation.
  3. (3)
    Proposes a challenging evaluation dataset Eval3DAIGC-198 with multimodal inputs, and proves the effectiveness of Idea23D through comprehensive user preference studies and 3D visual caption experiments.
@@ 2 Related Works
2.1 Self-refining Agents
@@
Our research builds on the self-refinement capability of large language models (LLMs). Recent studies show that LLMs are effective at self-refinement, when exploring unknown environments and tasks Xi et al. (2023); Madaan et al. (2024); Pan et al. (2023); Shinn et al. (2023); Lee et al. (2023). For example, projects such as Self-refineMadaan et al. (2024) and ReflexionShinn et al. (2023) utilize LLMs to iteratively critique their outputs and use feedback to improve predictions, resulting in significant performance improvements in natural language processing tasks. However, these approaches mainly excel in tasks dealing with natural language descriptions Shridhar et al. (2020). In contrast, our Idea23D project employs an iterative self-refinement system based on LMM in a multimodal environment, especially for interleaved inputs of text, images, and 3D models (IDEAs), in a different way other than the traditional approach focusing solely on natural language inputs.
2.2 Large Multimodal Model
@@
Building on the Large Language Model (LLM), the development of the Large Multimodal Model (LMM) marks an important evolution from unimodal to multimodal processing capabilities. Initial LLMs, such as the GPT family, focused on the generation of textual data, demonstrating superior capabilities in understanding and creating natural language Brown et al. (2020); Radford et al. (2019); OpenAI (2023d); Bai et al. (2023a, b); Zhang et al. (2023b); He et al. (2024). As the field evolves, the processing capabilities of the models expanded from pure text to include multimodal data including images, audio, and video Ramesh et al. (2021); Radford et al. (2021). For example, CLIP Radford et al. (2021) was the first to achieve cross-modal alignment between images and text, enabling cross-modal understanding between text and image. Some models Liu et al. (2024); Xu et al. (2024a); Chen et al. (2024b); Bai et al. (2023c); Wang et al. (2024a), demonstrates multimodal understanding and generation of mixed text, image, and video inputs. Further, projects such as Uni-3D Zhang et al. (2023d) and SDFusion Cheng et al. (2023) extend this concept to the 3D design domain, enabling good 3D model understanding, generation and reconstruction. LMM has also enabled open-set scene understanding in various settings Tian et al. (2023); Li et al. (2022, 2023); Liu et al. (2023e); Jin et al. (2023). Despite the progress made in the field of 3D understanding and generation, current LMMs can still only handle input content at the text and image level, and struggle to handle multimodal high-level inputs containing text, images, and 3D models (IDEAs in our case). In contrast, the Idea23D project takes a much larger step forward in multimodal 3D model generation. Our system is capable of processing not only single-modal inputs, but also composite multimodal inputs containing text, images, and 3D models at the same time.
2.3 Extensions of T-2-3D Models
@@
There is already a large literature extending seminal T-2-3D models, including variants enabling T-2-3D models to better follow user prompts Black et al. (2023); Chefer et al. (2023); Feng et al. (2022) , refine keywords in T-2-3D prompts for better visual quality Gu et al. (2023a), support for additional image inputs for image processing Brooks et al. (2023); Kawar et al. (2023), support for additional 3D model inputs for 3D model processing Cheng et al. (2023), 3D style migration Pan and Ke (2023); Ma et al. (2014); Segu et al. (2020), visual concept customization Wei et al. (2023); Kumari et al. (2023), using image generation models to generate 3D textures Chen et al. (2023a); Zeng et al. (2024); Chen et al. (2024a); Yu et al. (2024), and more. Going even further, Idea23D offers users a more natural way to design and create the 3D content they want. Similar to Visual ChatGPT Wu et al. (2023a), which extends ChatGPT’s OpenAI (2023e) ability to understand and generate 2D images, Idea23D is designed to provide a more unified and broadly applicable framework for automated 3D model design and generation. Idea23D extends the multimodal input and 3D model understanding and generation capabilities of the LMM OpenAI (2023d) using the T-2-I model Shi et al. (2020) and I-2-3D model Ramesh et al. (2021), as well as the multimodal input capabilities of the T-2-3D model Wang et al. (2024b); Poole et al. (2022); Lin et al. (2023); Wang et al. (2023); Metzer et al. (2023); Tsalicoglou et al. (2023); Qian et al. (2023); Liu et al. (2023d); Long et al. (2023) and other standard off-the-shelf algorithmic modules like multiview rendering.
3 Idea23D Framework
@@
The Idea23D framework represents a novel approach to generate detailed 3D models from high-level, abstract multimodal inputs (IDEAs), shown in Fig. 2. It integrates three LMM-based agents and several off-the-shelf tools for agents to invoke. Specifically, three agents are responsible for prompt generation (green box indexed by 1 in Fig. 2), model selection (green box indexed by 5 in Fig. 2) and feedback generation (green box indexed by 6 in Fig. 2). Two foundation models for T-2-I (purple box indexed by 2 in Fig. 2) and I-2-3D (purple box indexed by 3 in Fig. 2) are exploited together to turn natural language prompts into textured 3D models. As shown in Tab. 1, various foundation model variants are evaluated for comprehensiveness. A unique memory module enhances the system, retaining insights from previous iterations to optimize future outputs.
Figure 2: Overview of the framework of Idea23D, which employs LMM agents to unleash the T-2-3D model’s potential through iterative self-refinement to provide better T-2-3D prompts for the input user IDEA. Green rounded rectangles indicate steps completed by LMM agents. Purple rounded rectangles indicate T-2-3D modules, including T-2-I models and I-2-3D models. The yellow rounded rectangle indicates the off-the-shelf 3D model multi-view generation algorithm. The blue color indicates the memory module, which saves all the feedback from previous rounds, the best 3D model, and the best text prompt. Note that this cycle is fully automatically executed by LMM agents, without any human intervention.
@@ 3.2 Initial Prompt Generation
Here, $p_{\rm gen}$ represents a prompt facilitating the generation of T-2-3D prompts [^1] Note here $p_{\rm gen}$ and P are prompts for LMMs and T-2-3D models, respectively. All prompts are in our code.. This enables Idea23D to extract and interpret not only direct descriptions (e.g. $X^{\prime}$ with only $T$ is also acceptable) but also high-level concepts and intermixed modalities within the IDEA, such as images and 3D models.
@@
Idea23D transforms the set of N text prompts $\{P_{0},P_{1},\ldots,P_{N-1}\}$ into an equivalent set of N 3D models $\{D_{0},D_{1},\ldots,D_{N-1}\}$ utilizing T-2-3D models. As outlined in Sec. 3.2, T-2-3D encompasses a two-step conversion process: initial Text-to-Image (T-2-I) generation followed by Image-to-3D (I-2-3D) generation. To improve 3D model creation quality, we employ a background removal module  [^2] https://github.com/danielgatis/rembg on T-2-I outputs before I-2-3D processing. The T-2-3D function is: